40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Dec 4 19:34:22 2018, awade, DailyProgress, WOPO, Diabolo 532 nm power drop and retuning 18-12-04_DiaboloInside_annotate.jpegplot20181204_DiaboloSHGRaligned1064nm.pdfplot20181204_DiaboloSHGRErrSig.pdf20181204_RetuneDiaboloSHGAlignment.zip
    Reply  Thu Dec 6 20:04:50 2018, awade, DailyProgress, WOPO, Diabolo 532 nm power drop and retuning 
       Reply  Fri Dec 7 20:31:06 2018, awade, DailyProgress, WOPO, Diabolo: cleaning dust out of SHG unit, self heating issues 18-12-07_20-06-46_4774.jpg18-12-07_20-06-58_4776.jpg18-12-07_20-07-44_4780.jpgIMG_4785.JPGIMG_4766.MOV
          Reply  Thu Dec 13 14:01:13 2018, awade, DailyProgress, WOPO, Diabolo: contacted coherent 
             Reply  Thu Dec 13 16:24:06 2018, awade, DailyProgress, WOPO, Diabolo: baseline numbers for coherent plot20181218_PhaseMatchingCurveFastScanDiaboloSHG.pdf20181213_SHGScanWithLNTempTuning.zip
                Reply  Wed Dec 19 16:48:18 2018, awade, DailyProgress, WOPO, Diabolo: baseline numbers for coherent 
                   Reply  Thu Dec 20 19:20:50 2018, awade, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked plot20181220_SHGLockedOutputVsOvenSetpointTemperature.pdf20181220_SHGLockedPowerVsSetPointTemp.zip
                      Reply  Fri Dec 21 15:33:26 2018, Koji, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked 
                         Reply  Fri Dec 21 15:51:27 2018, awade, DailyProgress, WOPO, Diabolo: trialing SHG oven set temperature points to see if there is a maxima when locked 
                      Reply  Mon Dec 24 10:08:49 2018, awade, DailyProgress, WOPO, Diabolo: back to power with deliberate pre-misalignment 18-12-21_22-04-30_4812.jpg18-12-21_22-04-54_4813.jpg18-12-21_22-07-55_4814.jpg
                         Reply  Mon Dec 24 12:07:18 2018, Koji, DailyProgress, WOPO, Diabolo: back to power with deliberate pre-misalignment 
Message ID: 2280     Entry time: Thu Dec 13 14:01:13 2018     In reply to: 2277     Reply to this: 2281
Author: awade 
Type: DailyProgress 
Category: WOPO 
Subject: Diabolo: contacted coherent 

I've contacted the people at Coherent about getting instructions for repair or whether they can still service this unit as I don't want to waste a lot of futile time on trying to optimize a unit that might be broken/degraded.

The technicians I spoke to didn't seem to know a lot about the Diabolo unit but will forward on to their manufacture for further information/assistance.

 

Quote:

I've managed to vastly improve the power output of the SHG by doing a little cleaning inside the box.

After some consideration I decided that the only remaining possibility was something wrong inside the SHG. Its a hermetically sealed box so I cleaned around the area (to keep dust down) and then removed the lid.  Some of the screws were loose, so its likely that somebody has opened this before.  Below are a few pictures of the inside of the Diabolo SHG. 

 

 

  

Looking through an IR viewer I could see that there was a large absorber on the front face of the crystal.  I didn't get a picture but it was dead center and scattering a lot of light given the tightness of the beam. There was also other smaller scatters around the edge of the front surface. An clean N2 ionizing air gun was used to blow in and around the cavity and its electronics. There was a lot of dust blown out from inside the unit; the SHG has evidently been opened at some point in a dirty lab.  The spot in the middle of the front face of the SHG came loose with very gentle bushing with a corner of a lens cleaning tissue (MC-5).  I thoroughly flushed the inside of the unit again with 40 psi clean N2 and then replaced the lid.  Just before sealing the box I cracked the lid and flushed again with the N2 gun to fill the box with clean dry air.  The screws were sealed down with 2.25 Nm of torque.

I ramped the laser back on and after some small small alignment tweaks I was able to get an initial large boost in 532 nm laser output but it quickly dropped to the lower ~12 mW output again. It looks like, after lock, the output light goes through a few fringes and then either settles on a HOM or a weaker fundamental mode.  My guess is that in the locked state the crystal/mirror is heating up and mudslinging the cavity.  I don't really want to pull the unit apart into is constituent pieces to inspect.  I might need to have a closer look at the crystal front face absorber and the state of the input coupler mirror.  Over time some non-linear crystals do develop either grey tracking or permanent damage along the tighly focused points of 532 nm light.  It could be that the unit has gone past a tipping point but this is hard to see from just peaking in from the top of the unit.

Not sure what to do next.

---

Side note: Its seems very hard to make any alignment changes from the inside: my guess is that alignment is done from the three outside exposed hex screws on the front coupler. Its unlikely with such a short cavity much alignment changes would be needed.  For now its best to leave it as it is.

 

ELOG V3.1.3-